59. Uptown Industries just decided to save $3,000 a quarter for the next three years. The money
will earn 2.75 percent, compounded quarterly, and the first deposit will be made today. If the company
had wanted to deposit one lump sum today, rather than make quarterly deposits, how much would it
have had to deposit today to have the same amount saved at the end of the three years?
A. $34,441.56
B. $34,678.35
C. $33,428.87
D. $33,687.23
E. $34,998.01

- I believe the answer is B but I can not get this number on a financial calculator (BA 2).
Answer #1
PVAnnuity Due = c*((1-(1+ i)^(-n))/i)*(1 + i )
C = Cash flow per period
i = interest rate
n = number of payments
PV= 3000*((1-(1+ 2.75/400)^(-3*4))/(2.75/400))*(1+2.75/400)
PV = 34678.35

Set settings to BGN to calculate annuity due

N = 3*4

I/Y = 2.75/4

PMT = 3000

FV = 0

CPT PV